Two orthogonal Latin squares of order 10

A B C D E F G H I J
C A B E F G H I J D
B C A F G H I J D E
I H G J C D B E A F
J I H G D C E B F A
D J I A H E C F B G
E D J H A I F C G B
F E D B I A J G C H
G F E I B J A D H C
H G F C J B D A E I
a b c d e f g h i j
b c a j d e f g h i
c a b i j d e f g h
f g h e i a j b d c
g h i c f j a d b e
h i j f c g d a e b
i j d b g c h e a f
j d e g b h c i f a
d e f a h b i c j g
e f g h a i b j c d

Based on van Lint & Wilson, A Course in Combinatorics, chapter 22.